OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph. - 5 - 1 1 5

OpenStudy (anonymous):

its -1

OpenStudy (anonymous):

solve for y... y=-x+5 slope is -1

OpenStudy (anonymous):

first we write it in a form called slope-intercept form y=mx+b where (0,b) is the y-intercept of the line and m is the slope of the line. Let's do that: x+y=5 y=-x+5 now that we are in that form, we know that the coefficient of x will be the slope, so the slope is -1